# Crowd-SAM: SAM as a Smart Annotator for Object Detection in Crowded Scenes

### v1

@@ 3.4 Semantic-guided Mask Prediction
Given the proposals generated in  Sec. 3.3, this section aims to decode the dense prompts efficiently and discriminate the generated masks accurately. As depicted in  Fig. 3, each instance contains many prompts due to the density of grids. Supposing only one in-position prompt is required for mask prediction for SAM, decoding all the prompts would lead to not only a waste of computation but also cause more FPs due to some badly located prompts.
Figure 3: Illustration of EPS. The PWD-Net produce valid masks with a confidence threshold $T$. In each iteration, we prune prompts that fall inside valid masks (with a cross above).
Efficient Prompt Sampler. To address this challenge, we introduce the Efficient Prompt Sampler (EPS), which dynamically prunes prompts based on the confidence of decoded masks. This method is elaborated in  Algorithm 1. Beginning with the generated point prompts list $P_{G}$, EPS, in each iteration, samples a batch of prompts $P_{B}$ from $P_{G}$ using uniform random sampling. The sampled prompts $P_{B}$ are then appended to the output points list $P_{S}$. Subsequently, we employ SAM Generator with the batched prompts $P_{B}$ to produce masks $M$. Furthermore, we aggregate discriminant confidence scores $S$ from PWD-Net, which are employed to select valid masks $M^{\prime}$ using a score threshold $T$, by using $M^{\prime}=M[S>T]$. A detailed exposition of PWD-Net will be provided in the subsequent section. The valid masks represent regions that are believed to be well segmented, thus we remove the points that are already covered by any mask $m$ in $M^{\prime}$ from the prompts list $P_{G}$. The iteration halts when $P_{G}$ is empty. Furthermore, EPS establishes a stopping criterion with a parameter $K$, terminating the sampling process once the total sample count reaches this limit. This parameter is instrumental in managing the overall decoding cost.
Input: Feature extracted from image $I$ by image encoder of SAM: $F_{SAM}$; Generated prompt list: $P_{G}=\{p_{1},p_{2},...,p_{N}\}$; SAM Generator: $G$; Mask confidence threshold: $T$; Target sampled prompt list size: $K$;
Output: Sampled prompt list: $P_{S}$; Valid masks list: $M_{S}$.
@@ 11                   $P_{G}\leftarrow P_{G}\setminus\{p\}$
Part-Whole Discrimination Network. Given the raw masks predicted by SAM conditioned on the sampled batch of prompts $P_{B}$, we are inspired to design an automatic selector for choosing the best fitting mask. It should refine the outputs in two aspects: (i). help to refine the output IoU score according to the quality of related masks, if they are positive samples; (ii). help suppress the scores of samples that fall in the background regions. usion of those associated with the background.
Illustrated in  Fig. 2, for masks generated corresponding to $N$ prompts, we leverage the Mask Tokens and IoU Tokens within the mask decoder of SAM along with the sophisticated features extracted by the self-supervised pre-trained model DINOv2 [31]. ${\cal M}$ and ${\cal U}$ are responsible for mask decoding and IoU prediction in SAM Mask Decoder, respectively. Thus, we suppose that they contain shape-aware information, which is helpful in discriminating the mask.These components enable us to compute a discriminant confidence score $S$ for each specific prompt in a few-shot approach. Initially, the refined IoU score $S_{iou}$ is computed as follows:
@@
where $\text{Head}_{par}$ is a parallel IoU head, consisting of a MLP block, ${\cal U}\in\mathbb{R}^{N\times 1\times C}$ and ${\cal M}\in\mathbb{R}^{N\times 4\times C}$ denote the IoU Tokens and Mask Tokens respectively. Notice in  Eq. 2, the $S_{iou}$ is a sum of outputs from two individual heads, the parallel adapter head $\text{Head}_{par}$ and the original $\text{Head}_{IoU}$. We freeze the parameters of $\text{Head}_{IoU}$ to avoid overfitting in few-shot learning. Moreover, since these two tokens, ${\cal M}$ and ${\cal U}$ have different shapes, we repeat ${\cal U}$ by 4 times and then concat these two tokens in the channel dimension. The refined score $S_{iou}\in\mathbb{R}^{N\times 4}$ encapsulates the quality of the generated masks by assessing the texture-aware feature from ${\cal M}$ and ${\cal U}$.
Secondly, by harnessing the semantic feature embedded within the self-supervised pre-trained model DINOv2, along with the mask data $\hat{M}$, we calculate the discriminant score $S_{cls}$ :
@@
Here, we use ${\cal O}\in\mathbb{R}^{N\times C}$ to denote the extracted Semantic Token, $\hat{M}\in\mathbb{R}^{N\times H\times W\times 4}$ and $F_{DINO}\in\mathbb{R}^{N\times h\times w\times C}$ to represent the masks generated by SAM[29] and the features extracted by the DINOv2[31] model, respectively. We denote $d$ as a down-scale function that resizes the mask to $h\times w$ to be consistent with $F_{DINO}$. Pool is a global pooling function that conducts mean pooling on the x-axis and y-axis.The discriminant score $S_{cls}\in\mathbb{R}^{N\times 4}$ predicts whether masks belong to the foreground or background. The $\text{Head}_{CLS}$ shares the same parameters with the binary classifier introduced in  Sec. 3.3. Finally, we calculate a joint score of discrimination and estimate quality for masks by simply multiplying the two scores together: $S=S_{iou}\cdot S_{cls}$.
During training, prompts sampled within real masks are utilized as input. Regarding prompts within the foreground, the discriminant confidence score should aim to accurately predict the Intersection over Union (IoU) of generated and real masks. Conversely, for prompts within the background, the score should ideally output 00, indicating their association with the background. Hence, the loss function for this aspect is formulated as follows:
$s^{i}_{target}=\begin{cases}mIoU(m^{i},m^{i}_{GT}),&m^{i}_{GT}\in M^{bg}_{GT},\\ 0,&m^{i}_{GT}\in M^{fg}_{GT},\end{cases}$ (4)
@@ 3.5 Training and Inference
The total training loss of our framework is the combination of  Eq. 1 and Eq. 5:
@@
During inference, we select the mask with the highest confidence score $S$ among the $4$ masks as the output of PWD-Net, which we denote as $M\in\mathbb{R}^{N\times H\times W}$ and $S_{o}\in\mathbb{R}^{N}$. We adopt a window cropping strategy for better performance on small objects as [19]. This strategy slices the whole image into overlapping crops where each crop is processed individually. The final results are merged from the outputs of each crop. We apply non-maximum suppression on the final outputs to remove duplicate proposals.
@@ 4 Experiments
Table 2: Results on CrowdHuman[35] val set. All SAM-based methods adopt ViT-L[7] as the pre-trained backbone. SRCNN represents Sparse R-CNN[38] which is a baseline to [60]. * represents using the multi-crop trick.
Datasets. Following [60], we adopt CrowdHuman[35] as the benchmark to conduct main experiments and ablation studies. CrowdHuman[35] collects and annotates images containing crowded persons from the internet. It contains 15,000, 4,370, and 5,000 images for training, validation, and testing, respectively. We also test our method on CityPersons[55] for a realistic urban-scene evaluation. Additionally, to further investigate our method, we utilize OCC-Human [58] for evaluation, which is specially designed for occluded persons. For these pedestrian datasets, we utilize visible annotations (only including visible areas of an object) for training and evaluation. To validate the extensibility of Crowd-SAM, we also devise a multi-class version of Crowd-SAM by adding a multi-class classifier. We utilize 0.1 % percent of the COCO [23] trainval set for training and COCO val set for validation. We also validate our methods on an occluded subset of COCO, COCO-OCC [16] that is extracted by selecting images whose objects have a high overlapping ratio.
Implementation details. We utilize SAM (ViT-L) [19] and DINOv2 (ViT-L)[31] as the base models for all experiments. In the fine-tuning stage, their parameters are all frozen to avoid over-fitting. Instead of real GT, we use pseudo masks generated by SAM (ViT-L) as $S_{target}$ to supervise the learning of PWD-Net in Eq. 5. These generated pseudo labels are of high quality as shown in Fig. 4(c). We randomly pick points from the pseudo masks as positive training samples and points from the background part as negative training samples. In the training process, we use Adam[18] with a learning rate of $10^{-5}$, a weight decay of $10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$ for optimization. We train our module for 2000 iterations with a batch size of 1, which can be done on a single GTX 3090 Ti GPU in several minutes. For more details, please refer to the appendix.
Evaluation metrics. Following [60], we use AP with IoU threshold 0.5, MR-2, and Recall for our evaluation metrics. Generally, a higher AP, Recall, and lower MR-2 indicates a better performance.
@@ 4.1 Experimental Results on Pedestrian Detection
For a fair comparison, we re-implement the involved methods[56, 41, 60, 52, 32] with a 2×2$\times$2 × schedule using the visible annotations in the CrowdHuman [35] and CityPersons [55] dataset.
Main results. We compare Crowd-SAM with most related methods including fully supervised object detectors[41, 56, 52, 60], few shot object detectors[32] and SAM-based approaches[26, 19] for comprehensive understanding. Notice we use the visible annotation for which derives different results from those on full box annotation.
As shown in  Tab. 2, with only 10 labeled images, our method achieves comparable performance with full-supervised object detectors whose best result is 86.7% AP, achieved by DINO[52]. Particularly, Crowd-SAM outperforms an advanced anchor-free detector FCOS[41] by 2.1% AP. These results indicate that with proper adaptation techniques, SAM can achieve very competitive performance on complex pedestrian detection datasets like CrowdHuman. On the other hand, Crowd-SAM achieves SOTA performance on few-shot detection settings and outperforms all few-shot object detectors with a significant margin. DeFRCN [32] is a well-established few-shot object detector equipped with ResNet-101[13] that achieves 46.4% in AP. Notably, our method outperforms it by 32%, indicating the superiority of Crowd-SAM in the few-shot detection setting. The qualitative comparisons between ours and De-FRCN are depicted in  Fig. 4. For SAM-based approaches, our method largely leads the SAM baseline by 6% (with multi-crop) and 18.3% (w/o multi-crop) in the recall. Crowd-SAM is also superior to other methods such as Matcher[26].
Results on OccHuman and CityPersons. To investigate the performance of Crowd-SAM in occluded scenes, we compare it with an advanced supervised method Pose2Seg [58], and the results are shown in Tab. 3. It is noteworthy that Pose2Seg is a full supervised detector while Crowd-SAM is a few-shot detector. As can be seen from the results, Crowd-SAM leads Pose2Seg by 9.2% AP which demonstrates its robustness in occluded scenes. Also, we apply our method to a not-so-crowded but more realistic urban dataset, CityPersons [55], as reported in Tab. 4. Crowd-SAM outperforms TFA [42] by 2.7% AP, FSCE [37] by 1.1% AP, and De-FRCN [32] by 7.8% AP. In conclusion, our method remains competitive compared to advanced few-shot object detectors, even though it is not specifically designed for sparse scenes.
These results illustrate that our method has fully unleashed the power of vision foundation model models, i.e. SAM and DINO, in all of the crowded, occluded, and urban scenes.
@@
Mask R-CNN [12] R50-FPN 16.3 19.4 11.3
Pose2Seg [58] R50-FPN 22.2 26.1 15.0
Crowd-SAM ViT-L 31.4 26.5 17.7
Table 3: Comparative results on OccHuman [58] val. APM and APH represent AP in moderate and hard cases according to occlusion ratio, respectively.
@@
Table 4: Comparative results on CityPersons[55] val.
@@
Faster R-CNN [34] R-50-FPN 29.7 50.0 33.5 53.7
BCNet [16] R-50-FPN 31.7 51.1 34.6 54.4
Crowd-SAM ViT-L 20.6 31.5 22.0 33.7
Table 5: Comparative results on COCO [23] and COCO-OCC [16] val.
@@ 4.2 Experimental Results on Multi-class Object Detection
To further explore the extensibility of our methods in a more popular setting, we devise a multi-class Crowd-SAM. The multi-class version of Crowd-SAM is slightly different by replacing the binary classifier with a multi-class classifier. We validate multi-class Crowd-SAM on COCO [23], a widely adopted common object detection benchmark, and COCO-OCC [16] which is a split of COCO that mainly consists of images with high occlusion ratio.
We compare our method with two supervised detectors, Faster R-CNN [34] and BCNet [16], and report the results in Tab. 5. It is seen that our Crowd-SAM is comparable to supervised detectors in both datasets and drops only 1.4 APwhen comparing COCO and COCO-OCC. This minor drop indicates that our method is robust to the occlusion.
@@
Table 6: Ablation studies on the main components in Crowd-SAM. FG location represents the use of binary classifier on DINOv2 features. The last line represents a SAM baseline with a standard $32\times 32$ grid as inputs.

### v2

@@ 3.4 Semantic-guided Mask Prediction
Given the proposals generated in Sec. 3.3, our further aim is to efficiently decode the dense prompts and accurately discriminate the generated masks. As depicted in Fig. 3, each instance contains a set of prompts due to the density of grids. Supposing that only one in-position prompt is required for mask prediction in SAM, decoding all the prompts would lead to not only a waste of computation but also more FPs for some poorly located prompts.
Figure 3: Illustration of EPS. PWD-Net produces valid masks with a threshold. In each iteration, we prune prompts (with a cross above) that fall inside valid masks .
Efficient Prompt Sampler (EPS). To address this challenge, we introduce EPS, which dynamically prunes prompts based on the confidence of decoded masks. This method is elaborated in Algorithm 1. Beginning with the generated point prompts list $P_{G}$, EPS, in each iteration, samples a batch of prompts $P_{B}$ from $P_{G}$ using uniform random sampling. The sampled prompts $P_{B}$ are then appended to the output points list $P_{S}$. Subsequently, we employ the SAM generator with the batched prompts $P_{B}$ to produce masks $M$. Furthermore, we aggregate discriminant confidence scores $S$ from PWD-Net, which are employed to select valid masks $M^{\prime}$ using a score threshold $T$, by using $M^{\prime}=M[S>T]$. Refer to a detailed exposition of PWD-Net in the subsequent section. The valid masks represent regions that are believed to be well segmented, and we thus remove the points that have already been covered by any mask $m$ in $M^{\prime}$ from the prompts list $P_{G}$. The iteration halts when $P_{G}$ is empty. Additionally, EPS establishes a stopping criterion with a parameter $K$, terminating the sampling process once the total sample count reaches this limit. This parameter is instrumental in managing the overall decoding cost. In our experiments, $K$ is empirically set to 500 to strike a balance between efficiency and recall.
Input: feature extracted from image $I$ by image encoder of SAM: $F_{SAM}$; generated prompt list: $P_{G}=\{p_{1},p_{2},...,p_{N}\}$; SAM generator: $G$; mask confidence threshold: $T$; target sampled prompt list size: $K$
Output: sampled prompt list: $P_{S}$; valid mask list: $M_{S}$
@@ 11                   $P_{G}\leftarrow P_{G}\setminus\{p\}$
Part-Whole Discrimination Network (PWD-Net). Given the raw masks predicted by SAM conditioned on the sampled batch of prompts $P_{B}$, we design an automatic selector for choosing the best-fitting mask. It is expected to optimize the outputs in two aspects: (i) refining the output IoU score according to the quality of related masks if they are positive samples and (ii) suppressing the scores of samples that fall in background regions.
Illustrated in Fig. 2, for the masks generated corresponding to $N$ prompts, we leverage the Mask Tokens and IoU Tokens within the mask decoder of SAM along with the sophisticated features extracted by the self-supervised pre-trained model DINOv2 [33]. ${\cal M}$ and ${\cal U}$ are responsible for mask decoding and IoU prediction in the SAM Mask Decoder, respectively. Thus, we suppose that they contain shape-aware information, which is helpful in discriminating the mask. These components enable us to compute a discriminant confidence score $S$ for each specific prompt in a few-shot scenario. Initially, the refined IoU score $S_{iou}$ is computed as follows:
@@
where $\text{Head}_{par}$ is a parallel IoU head, consisting of an MLP block; ${\cal U}\in\mathbb{R}^{N\times 1\times C}$ and ${\cal M}\in\mathbb{R}^{N\times 4\times C}$ denote the IoU Tokens and Mask Tokens respectively. Notice in Eq. 2, $S_{iou}$ is a sum of outputs from two individual heads, the parallel adapter head $\text{Head}_{par}$ and the original $\text{Head}_{IoU}$. We freeze the parameters of $\text{Head}_{IoU}$ to avoid overfitting in few-shot learning. Moreover, since these two tokens, ${\cal M}$ and ${\cal U}$ have different shapes, we repeat ${\cal U}$ by 4 times and concatenate these two tokens in the channel dimension. The refined score $S_{iou}\in\mathbb{R}^{N\times 4}$ encapsulates the quality of the generated masks by assessing the texture-aware feature from ${\cal M}$ and ${\cal U}$.
Further, by harnessing the semantic feature embedded in the self-supervised pre-trained model DINOv2, along with the mask data $\hat{M}$, we calculate the discriminant score $S_{cls}$ :
@@
Here, ${\cal O}\in\mathbb{R}^{N\times C}$ denotes the extracted Semantic Token, and $\hat{M}\in\mathbb{R}^{N\times H\times W\times 4}$ and $F_{DINO}\in\mathbb{R}^{N\times h\times w\times C}$ represent the masks generated by SAM[31] and the features extracted by DINOv2, respectively. We denote $d$ as a down-scale function that resizes the mask to $h\times w$, consistent with $F_{DINO}$. Pool is a global pooling function that conducts mean pooling on the x-axis and y-axis and $\circ$ is the Hadamard product. The discriminant score $S_{cls}\in\mathbb{R}^{N\times 4}$ predicts whether masks belong to the foreground or background. $\text{Head}_{CLS}$ shares the same parameters with the binary classifier introduced in Sec. 3.3. Finally, we calculate a joint score of discrimination and estimate the quality for masks by simply multiplying the two scores: $S=S_{iou}\cdot S_{cls}$.
During training, prompts sampled from real masks are taken as input. Regarding the prompts within the foreground, the discriminant confidence score aims to accurately predict the IoU of the generated and real masks. Conversely, for the prompts within the background, the score is ideally 00. Hence, the loss function for this aspect is formulated as follows:
$s^{i}_{target}=\begin{cases}IoU(m^{i},m^{i}_{GT}),&m^{i}_{GT}\in M^{bg}_{GT},\\ 0,&m^{i}_{GT}\in M^{fg}_{GT},\end{cases}$ (4)
@@ 3.5 Training and Inference
The total training loss of the entire framework combines Eq. 1 and Eq. 5:
@@
At inference, we select the mask with the highest confidence score $S$ among the $4$ masks as the output of PWD-Net, which we denote as $M\in\mathbb{R}^{N\times H\times W}$ and $S_{o}\in\mathbb{R}^{N}$. We adopt a window cropping strategy to enhance the performance on small objects as [20]. This strategy slices the whole image into overlapping crops where each crop is individually processed. The final results are merged from the outputs of each crop and here we apply NMS to remove duplicate proposals.
@@ 4 Experiments
Table 2: Comparative results (%) on CrowdHuman [37] val. All the SAM-based methods adopt ViT-L [7] as the pre-trained backbone (SRCNN denotes Sparse R-CNN [40], a baseline in [62] and * represents using the multi-crop trick).
@@
Datasets. Following [62], we adopt CrowdHuman[37] as the benchmark to conduct main experiments and ablation studies. CrowdHuman [37] collects and annotates images containing crowded persons from the Internet. It contains 15,000, 4,370, and 5,000 images for training, validation, and testing, respectively. We also evaluate our method on CityPersons[58] for a realistic urban-scene scenario. Additionally, we utilize OCC-Human [61], which is specially reputed for occluded persons. For these pedestrian datasets, we use visible annotations (only including visible areas of an object) for training and evaluation. To validate the extensibility of Crowd-SAM, we further devise a multi-class version of Crowd-SAM by adding a multi-class classifier. We employ 0.1 % percent of the COCO [24] trainval set for training and the COCO val set for validation. Besides, we validate our method on a subset with occluded objects on COCO, i.e. COCO-OCC [17], extracted by selecting the images whose objects have a high overlapping ratio.
Implementation Details. We utilize SAM (ViT-L) [20] and DINOv2 (ViT-L) [33] as the base models for all experiments. In the fine-tuning stage, all their parameters are frozen to avoid over-fitting. Instead of real GT, we use the pseudo masks generated by SAM as $S_{target}$ to supervise the learning of PWD-Net in Eq. 5. These generated pseudo labels are of high quality as shown in Fig. 4(c). We randomly pick the points from the pseudo masks as positive training samples and the ones from the background as negative training samples. In the training process, we use Adam[19] with a learning rate of $10^{-5}$, a weight decay of $10^{-4}$, $\beta_{1}=0.9$, and $\beta_{2}=0.99$ for optimization. We train our module for 2,000 iterations with a batch size of 1, which can be done on a single GTX 3090 Ti GPU in several minutes. For more details, please refer to the Appendix.
Evaluation Metrics. Following [62], we use AP with IoU threshold at 0.5, MR-2, and Recall as metrics. Generally, a higher AP, Recall, and lower MR-2 value indicates better performance.
@@ 4.1 Experimental Results on Pedestrian Detection
For fair comparison, we re-implement the counterparts[59, 43, 62, 55, 34] with a 2×2$\times$2 × schedule using visible annotations in the CrowdHuman [37] and CityPersons [58] datasets.
Main Results. We compare Crowd-SAM with most related methods including fully-supervised object detectors [43, 59, 55, 62], few shot object detectors [34] and SAM-based methods [28, 20]. Notice that we use visible annotations for which derive different results from those on full box annotations.
As Tab. 2 shows, with only 10 labeled images, Crowd-SAM achieves comparable performance with full-supervised object detectors whose best result is 86.7% AP, delivered by DINO[55]. Particularly, Crowd-SAM outperforms an advanced anchor-free detector FCOS[43] by 2.1% AP. These results indicate that by using proper adaptation techniques, SAM can reach very competitive performance on complex pedestrian detection datasets like CrowdHuman. On the other side, Crowd-SAM achieves SOTA performance on few-shot detection settings and outperforms all the few-shot object detectors by a significant margin. DeFRCN [34] is a well-established few-shot object detector equipped with ResNet-101 [14] that reports 46.4% in AP. Notably, our method exceeds it by 32%, indicating the superiority in the few-shot detection setting. The qualitative comparison between Crowd-SAM and De-FRCN is depicted in Fig. 4. For SAM-based methods, our Crowd-SAM largely leads the SAM baseline by 6% (with multi-crop) and 18.3% (w/o multi-crop) in Recall. Besides, Crowd-SAM is also superior to the other methods such as Matcher [28].
Results on OccHuman and CityPersons. To investigate the performance of Crowd-SAM in occluded scenes, we compare it with an advanced counterpart Pose2Seg [61], and the results are shown in Tab. 3. It is noteworthy that Pose2Seg is a fully-supervised detector while Crowd-SAM is a few-shot detector. As can be seen from the results, Crowd-SAM leads Pose2Seg by 9.2% AP which demonstrates its robustness in occluded scenes. Also, we apply our method to a not-so-crowded but more realistic urban dataset, i.e. CityPersons [58], as reported in Tab. 4. Crowd-SAM outperforms TFA [44] by 2.7% AP, FSCE [39] by 1.1% AP, and De-FRCN [34] by 7.8% AP. In conclusion, our method remains competitive compared to advanced few-shot object detectors, even though it is not specifically designed for sparse scenes.
These results illustrate that Crowd-SAM well unleashes the power of vision foundation models, i.e. SAM and DINO, in all of the crowded, occluded, and urban scenes.
Table 3: Comparative results (%) on OccHuman val, where APM and APH represent AP in moderate and hard cases according to occlusion ratios, respectively.
@@
Mask R-CNN [13] ResNet50-FPN [14] 16.3 19.4 11.3
Pose2Seg [61] ResNet50-FPN [14] 22.2 26.1 15.0
Crowd-SAM ViT-L [7] 31.4 26.5 17.7
Table 4: Comparative results (%) on CityPersons val.
@@
Table 5: Comparative results (%) on COCO val and COCO-OCC.
@@
Faster R-CNN [36] ResNet50-FPN [14] 29.7 50.0 33.5 53.7
BCNet [17] ResNet50-FPN [14] 31.7 51.1 34.6 54.4
Crowd-SAM ViT-L [7] 20.6 31.5 22.0 33.7
@@ 4.2 Experimental Results on Multi-class Object Detection
To further explore the extensibility in a more popular setting, we devise a multi-class Crowd-SAM. The multi-class version of Crowd-SAM is slightly different by replacing the binary classifier with a multi-class one. We then validate multi-class Crowd-SAM on COCO [24], a widely adopted common object detection benchmark, and COCO-OCC [17] which is a split of COCO that mainly consists of images with a high occlusion ratio.
We compare our method with two supervised detectors, i.e. Faster R-CNN [36] and BCNet [17], and report the results in Tab. 5. It can be seen that our Crowd-SAM is comparable to the supervised detectors on both datasets and drops only 1.4 APwhen comparing those of COCO and COCO-OCC. This minor drop indicates that our method is robust to occlusions.
Table 6: Ablation results (%) on the main components in Crowd-SAM. FG location represents the use of the binary classifier on DINOv2 features. The last line represents a SAM baseline with a standard $32\times 32$ grid as inputs.